Jerome will be buying a used car for ​$7,000 in 3 years. How much money should he ask his parents for now so​ that, if he invests it at 8​% compounded
continuously​, he will have enough to buy the​ car?
Jerome should ask for $__

Answer by Boreal(15235)

x*e^(0.08*3)=7000
x*e^0.024=7000
x=7000/1.271 (round at end)
=$5506.40